OpenStudy (anonymous):

A branch falls from a tree. How fast is the branch moving after 0.28 seconds?

OpenStudy (theeric):

This is an issue where we can use the definition of acceleration. That is, the change in velocity per some change in time. Also, acceleration is gravitational acceleration. \(a=\dfrac{\Delta v}{\Delta t}\) And \(g=a\), right? So \(g=\dfrac{\Delta v}{\Delta t}\) So, we have that the initial velocity is 0 at the start of the fall. So, you can think that the change in velocity is the new velocity. If you want to see the math that agrees, \(\Delta v=v_f-v_0=v_f-0=v_f\) Does that help?